OpenStudy (crashonce):

express sin(4A)/sinA as a function of cos A

OpenStudy (junyang96):

use the double angle formulas \[\sin(2x)=2\sin(x)\cos(x)\]\[\cos(2x)=2cos ^{2}(x)-1\]